OpenStudy (anonymous):

Solve the system of equations. Enter your answer as an ordered pair. 6x + 7y = 96 6x - 7y = 12

OpenStudy (anonymous):

by adding u will get, 12x = 108 x = 108/12 x = 9 6(9)-7y = 12 54-7y = 12 -7y = 12-54 -7y = -42 y = -42/-7 y = 6

OpenStudy (anonymous):

u r in 10th grade.right

OpenStudy (anonymous):

mhm